OpenStudy (anonymous):

\[Evaluate \int\limits \frac{x ^{2}}{x ^{2}+1}\]

OpenStudy (anonymous):

Try to solve the numerator as: \(x^2 + 1 - 1\) So, It will become: \[\int\limits_{}^{}\frac{x^2 + 1-1}{x^2 + 1}dx = \int\limits_{}^{}\frac{x^2 + 1}{x^2 +1}dx - \int\limits \frac{1}{x^2 + 1}dx \implies \int\limits dx - \int\limits \frac{1}{x^2 + 1}dx\]

OpenStudy (anonymous):

Try to solve this..

OpenStudy (anonymous):

ok thanks. :)